OpenStudy (anonymous):

Mario has a pizzeria. The times to prepare and deliver pizzas are normally distributed with mean 20 minutes and standard deviation 5 minutes. Use the empirical rule to find an interval in which about 95% of the times will lie. A) (15 minutes, 25 minutes) B) (10 minutes, 30 minutes) C) (5 minutes, 35 minutes) D) (0 minutes, 40 minutes)

OpenStudy (amistre64):

whats the sd for 95%? 1 2 or 3 sds?

OpenStudy (amistre64):

if you can recall the empirical rules, then the question basically answers itself :)

OpenStudy (kropot72):

The empirical rule says that about 95% of the area under the nornal distribution curve is included within a distance of about +- 2 standard deviations from the mean. In this case 2 standard deviations = 2 * 5 = 10 minutes. The lower limit of the interval = 20 - 10 minutes The upper limit of the interval = 20 + 10 minutes. Can you now finish it?